OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. -x+5y=14(-5,-2)

OpenStudy (anonymous):

options A- y=-5x-27 B- y= -1/5x-27 C- y=-1/5x-15 D- y= 5x-27

OpenStudy (anonymous):

To start with you should get the original equation in Point-Slope form ( Y=mX+b form) Then the slope of a perpendicular line is 1/m

OpenStudy (anonymous):

-1/m *